fertilizers contain ammonium ion which can be toxic Suppose the toxicity level of the ammonium ion is 0.38 mol / L and above If 8.19 g of ammonium phosphate are dissolve in water to create 432 Ml of solution would this be considered toxic
1
Expert's answer
2020-02-28T12:14:49-0500

The dissociation of ammonium phosphate in water can be described as following:

(NH4)3PO4 = 3NH4+ + PO4-

From here, the mass of ammonia ions in the solution equals:

m (NH4+) = [m((NH4)3PO4) × 3 × Mr(NH4+) ]/Mr((NH4)3PO4) = (8.19 g × 3 × 18.04 g/mol) / 149.09 g/mol = 2.97 g

The molarity of ammonium ions in the solution equals:

n (NH4+) = m(NH4+) / [Mr(NH4+) × V(solution)] = 2.97 g / (18.04 g/mol × 0.432 L) = 0.381 mol/L


Answer: As toxicity the toxicity level of the ammonium ion is 0.38 mol / L, the solution of ammonium phosphate is toxic.